# Static Three-Dimensional Structures Determine Fast Dynamics Between Distal Loci Pairs in Interphase Chromosomes

### v2

@@ II Results
Next, we calculated the two-point mean square displacement $M_{2}(t)$ using Eq. 3 and $M_{2}(t)=2\langle r_{ij}^{2}\rangle-2G_{2}^{ij}(t)$. Use of Eq. 3 requires knowledge of the connectivity matrix $\bm{K}$, which is the byproduct of the determination of the 3D chromatin coordinates (Eq. 2). Fig. 2(d) shows the time dependence of $M_{2}(t)$ predicted by our theory for the pairs of E-P distances shown in Fig. 2(b). At long times $M_{2}(t)$ saturates, approaching the different equilibrium values that depend on the given pair. The rate of approach depends on the specific enhancer and promoter pair.
@@
We then calculated the relaxation time $\tau$. Because the absolute value of the friction coefficient is unknown, we tuned it to achieve the best agreement between the theoretical prediction and experimental value for $\tau$. The fit parameter yields the unit length $l_{0}=147\mathrm{nm}$ and unit time is $\tau_{0}\approx 3.1\mathrm{s}$. We used these values to calculate the theoretical predictions for $\tau$ versus genomic distance $s$ to compare with experiments. The theoretical prediction for the scaling exponent, $\gamma$ in $\tau\sim s^{\gamma}$, is $\approx 0.8$ and for experimental data is $\approx 0.7$ (Fig. 2(e)). It is important to note that while $l_{0}$ and $\tau_{0}$ are adjustable parameters used to calculate the absolute values of $\tau$, they do not affect the scaling exponent $\alpha$. We consider the agreement for $\tau$, and especially $\gamma$, between experiment and theory to be striking because the only information that is used in the calculation is the Micro-C static contact map.
@@
Randomly Shuffled Sequence: We then wondered if the introduction of randomness in the Micro-C contact map would lead to a discrepancy between theory and experiment. To this end, we randomly shuffled the pairwise distances in the distance map but retained the first off-diagonal elements. In this way, the polymeric nature of the structure is preserved but the specific WT (wild type) pattern in the Micro-C contact map is destroyed. We then applied HIPPS-DIMES on the shuffled distance map to obtain the connectivity matrix. Comparison of the contact maps between the WT and the randomly shuffled sequence shows (Supplementary Figure 6(a)) that the pattern in the WT contact map is fully disrupted. Supplementary Figure 6(b) shows the mean pairwise distances as the function of the genomic distance. For $s<10^{2}$ bps, the mean distance grows roughly as a power law with an exponent of 0.6, demonstrating the polymeric nature is preserved in the random shuffled system. At $s>10^{2}$ bps, the mean distances reach the plateau, which is a result of random shuffling. The relaxation time $\tau$ for the shuffled sequence is insensitive to the genomic distance (Supplementary Figure 6(c)), which is consistent with the saturation of mean distances. The purpose of this calculation is to show that the scaling of $\tau$ as a function of $s$ cannot be captured in random copolymers. The chromatin sequence, reflecting the patterns of activity depicted in histone modifications, and the associated 3D structures should be accounted for precisely.
@@
Figure 3: (a) Comparison between the mean spatial distance matrix (lower triangle) and the map for relaxation time $\tau$ (upper triangle). (b) Scatter plot of relaxation time $\tau$ versus the corresponding spatial distances $\langle r\rangle$. The black marks are the experimental measurements. The blue markers are the theoretical calculations. The Red dashed line is the fit to the theoretical results.
@@
Loci-dependent relaxation times: Given the remarkable success of our theoretical approach in quantitatively explaining the experimental findings, we calculated all the pairwise relaxation time $\tau_{ij}$ where $i,j$ are the loci pair indices. The value of $\tau_{ij}$ depends on both $i$ and $j$, and not merely on the genomic distance $s=|i-j|$ as in the case of homopolymer. On an average, the relaxation time correlates with both $s$ and the mean spatial distance $\langle r_{ij}\rangle$ in non-trivial manner. Fig. 3(a) shows the mean spatial distance map and the relaxation time map, clearly establishing the correlation between the two quantities. Fig. 3(b) shows the scatter plot of $\langle r_{ij}\rangle$ versus $\tau_{ij}$. The results show that, on an average, they are related as $\tau_{ij}\sim\langle r_{ij}\rangle^{2.6}$, which differs substantially from the prediction for the Rouse and the FG models. In particular, the scaling exponent $\sim 2.6$ is significantly smaller than the Rouse model prediction ($=4$) and the value for FG ($=5$).
@@
We then wondered whether the observed scaling can be deduced by considering an effective homopolymer model, in which the mean distance matrix is calculated as the average of the wild-type (WT) distance map over fixed genomic distances. The calculation is intended to assess if a modified scaling relation could be used with the mean distance that is calculated from the contact maps. We computed $r(s)$ by averaging $\langle r_{ij}\rangle$ over $s=|j-i|$, and then applied HIPPS-DIMES to obtain the relaxation time. Supplementary Figure 7 shows the calculated average distance map, demonstrating that $r(s)\sim s^{1/4}$ for $s>10^{2}\ \mathrm{kb}$. Supplementary Figures 7(c) and 7(d) show that the relaxation time scales with genomic distance as $\tau\sim s^{1.1}$ and with mean pairwise distance as $\tau\sim r^{4}$. Both of these scaling relations are different from the results obtained by considering the full WT contact map. This further demonstrates that arguments in standard polymer physics do not hold for the chromatin, thus underscoring the importance of considering the complete information embedded in the WT contact map. Together, these results show that the relaxation process between a pair of chromatin loci is much faster than predicted by standard polymer models, which provides a structural basis for interpreting the key experimental finding [36]. Importantly, the rapid dynamics between distal loci can be explained by taking into account the actual 3D coordinates of the chromosomes.
@@
Finally, we examined whether promoter-enhancer relaxation is faster than that of non-promoter-enhancer pairs. For this purpose, we computed the distribution of relaxation times for all loci pairs with the same genomic separation as the corresponding enhancer-promoter pairs. The results, shown in Supplementary Figure 8, reveal distinct trends based on the genomic distance. For very short genomic distances (enhancer-promoter pairs 1 and 2, with separations of 58 kb and 82 kb), the mean relaxation times (133 s and 134 s, respectively) are significantly shorter than the mean relaxation times of all pairs at the same genomic distance (207 s and 240 s). In contrast, for larger genomic distances (all other enhancer-promoter pairs), the mean relaxation times are consistently longer than the corresponding mean values of all pairs with the same separation. These observations suggest that the relaxation dynamics of enhancer-promoter pairs may be influenced by genomic distance in a manner distinct from general locus pairs.
@@
Plausible mechanism for rapid relaxation dynamics in chromatin: To explore the underlying mechanism for the rapid relaxation times between chromatin loci, we calculated the spectrum of eigenvalues, $\lambda_{p}$, associated with, $\bm{K}$, the connectivity matrix. Interestingly, Fig. 4(a) shows that the scaling of $|\lambda_{p}|$ with $p$ has a complex structure. There are three distinct regimes in the variation of $\lambda_{p}$ with the normal mode $p$. For $p\leq 10$, we find that $|\lambda_{p}|\approx p^{1.2}$. In the second regime, $10\leq p\leq 50$ the eigenvalues increase as $|\lambda_{p}|\sim p^{3}$. Finally, $|\lambda_{p}|\sim p^{1.5}$ for $p\geq 50$. The complicated spectrum for chromatin should be contrasted with the Rouse model for which $|\lambda_{p}|\sim p^{2}$ where the inverse of $|\lambda_{p}|$ maybe interpreted as the relaxation time of $N/p$ segments of the chain. Notably, the decrease in the scaling exponents in the small $p$ regime (Fig. 4(a)) compared to the $p^{2}$ scaling of the Rouse model supports the finding that chromatin relaxes more rapidly, consistent with the results for relaxation time $\tau$. In general, for a homopolymer, we expect that $|\lambda_{p}|\sim p^{x}$, which implies that the relaxation times scale with respect to the chain length with the same exponent, $\tau_{ee}\sim N^{x}$. For instance, for the Rouse model, $|\lambda_{p}|\sim p^{2}$ and $\tau_{ee}\sim N^{2}$. If we assume that a similar power law relationship holds between $\tau_{ee}$ and $p$ in chromosomes, then expect that $\tau_{ee}\sim N^{1.2}$ if $s\gtrapprox$ 400 kb. To test this prediction, we calculated the end-to-end-distance relaxation times in chromosomes with different lengths. The HIPPS-DIMES-based calculation shows that the end-to-end relaxation time roughly scales linearly with $N$ (Fig. 4(b)) as $\tau_{ee}$ varies by over five orders of magnitude. We also computed the eigenvalues for the randomly shuffled system. Supplementary Figure 9(a) shows that $|\lambda_{p}|$ becomes independent of $p$ for $p\leq 40$, which is is consistent with the findings that $\tau$ is insensitive to the genomic distance in the shuffled system. Together, these results suggest that the dynamics of chromosomes are dependent on the sequence and the length scale, which is reflected in the observation that the $|\lambda_{p}|$ exhibits three distinct scaling regimes at different $p$ (different length scales).
@@
Figure 4: (a) Scaling of the eigvenvalues, $\lambda_{p}$, of the connectivity matris as a function of $p$. Three distinct scaling regimes are shown. (b) End-to-end relaxation time $\tau$ as a function of chain length $N$. The dashed line is the fit.
@@
First passage time of contact between distal loci: A functionally relevant biophysical property related to the two-point relaxation time is the first passage time of contact between a pair of chromatin loci. A simpler, well-studied, and instructive version of this problem in polymer physics is the cyclization process, which concerns the first passage time for two ends of a polymer chain to meet [48, 49, 50, 46]. Let us denote the first passage time of contact as $\tau_{c}$, which is determined by the search process by which two loci meet. It can be shown that $\tau_{c}$ is directly connected to the two-point relaxation and is governed by the relaxation dynamics between the loci [46, 50], assuming that the threshold for establishing contact is not small. Using a contact threshold of $r_{c}=147\mathrm{nm}$, we estimated the mean FPT, $\langle\tau_{c}\rangle$, for all pairs of the chromatin loci. Fig. 5(a) shows that the domain along the diagonal in the $\tau_{c}$ map visually matches the contact domains in the contact map. We then calculated $\tau_{c}$ by averaging over the spatial distance $r$ with fixed genomic distance $s$. Fig. 5(b) shows that, on an average, the mean $\tau_{c}$ of contact between two chromatin loci scales as $\langle r\rangle^{3.4}$.
@@
We also compared our theoretical predictions with results computed using experimental loci trajectory data. The method for calculating the mean FPT from experimental trajectories is described in the Supplementary Information. As shown in Fig. 5(b), the experimental results (represented by triangle symbols) align closely with our predictions, demonstrating excellent quantitative agreement between the experiment and the theory.
@@
Interestingly, the calculated scaling exponent of 3.4 is close to 3, which is the theoretical prediction by Szabo, Schulten, and Schulten (SSS) [49] who derived the first passage time of contact under the assumption that single-particle diffusion is governed by dynamics in a potential of mean force, which can be calculated analytically for the Rouse model. The potential of mean force is determined from the distribution of distances between two loci which further leads to [49, 46, 51],
@@
where $p(x)$ is the equilibrium distribution of distances $x$ between the two loci with the mean distance $\langle r\rangle$, $r_{c}$ is the threshold distance distance for contact, and $D$ is the effective diffusion constant. It can be shown using Eq. 4 that $\tau_{c,\mathrm{SSS}}\sim D^{-1}r_{c}^{-1}\langle r\rangle^{3}$. This result is consistent with the visual similarity between the contact map and the contact time map shown in Fig. 5(a), as the contact probability scales with the mean distance with an exponent of 3 (in the HIPPS-DIMES model). It is intriguing that the SSS theory, developed for contact formation between a single pair, qualitatively captures the establishment of contacts involving multiple loci pairs. In contrast, the Rouse model does not predict the correct scaling of $\tau_{c}$ as a function of $\langle r\rangle$.
@@
The value of $\tau_{c}$ is important in describing the dynamics of promoter-enhancer communication. If we assume that gene expression is initiated by the formation of contact between promoter and enhancer then the transcription rate can be expressed as $k=1/(\tau_{\mathrm{d}}+\tau_{c})$, where $\tau_{\mathrm{d}}$ is the time required for the downstream processes that ensue after the establishment of contact. By expressing $\tau_{c}$ as a function of the contact probability $p_{c}$, with $\tau_{c}\sim\tau_{0}p_{c}^{-\theta}$, we obtain $\hat{k}=k/k_{\mathrm{max}}=1/(1+(\tau_{0}/\tau_{d})p_{c}^{-\theta})$. This equation is the dynamic analog of the Hill equation with the “cooperativity” parameter $\theta$. Such an equation has been used to model the mean mRNA number as a function of the contact probability between promoter and enhancer [52].
@@
Figure 5: (a) Comparison between heatmap for the mean first passage time of contact between chromatin loci (upper triangle), $\langle\tau_{c}\rangle$, and the contact map (lower triangle). The contact is defined with a distance threshold of $r_{c}=147\mathrm{nm}$. (b) $\langle\tau_{c}\rangle$ scales as the mean distance $\langle r\rangle$ as a power-law. The dashed line is a guide for the eye. Triangle symbols are the data computed using experimental trajectories data [36].
@@
Single-Loci Dynamics and Centrality Measure: Having demonstrated that our theory quantitatively reproduces the experimental data for the two-point dynamics in chromatin loci, we explore the predictions of single-locus dynamics. The single-locus mean square displacement, $M_{1}(t)$, is computed using Eq. 14 in the Supplementary Information. The prediction for $M_{1}(t)$ is shown in Fig. 6(a), where each line represents a single chromatin locus. Between the time scale of $1\mathrm{\ s}<t<10^{5}\mathrm{\ s}$, $M_{1}(t)$ scales approximately with an exponent of $\sim 0.5$, which is close to the prediction of the Rouse model. Next, we calculated the diffusion exponent $\alpha$ and diffusion constant $D$ for each locus by fitting $M_{1}(t)$ in the time range $1\mathrm{s}<t<10^{5}\mathrm{s}$ using $M_{1}(t)=Dt^{\alpha}$. The histogram of $\alpha$ and $D$ reported in Fig. 6(b), yields the ensemble averages of $\langle\alpha\rangle=0.52$ and $\langle D\rangle=0.024\ \mu\mathrm{m^{2}s^{-1/2}}$. The distribution of $\alpha$ is broad as is the spread in the effective diffusion coefficient (Fig. 6(b). Our results show that single-locus diffusion is heterogeneous.
@@
We then investigated the factors that determine the variance in single-locus diffusion. We hypothesize that chromatin loci should generally diffuse more slowly if they have higher local density. Inspired by concepts in graph theory [53], we defined the closeness centrality measure based on the mean pairwise distances. Let us define the centrality of a single locus as $C_{i}$, where $i$ is the locus index, as, $C_{i}=\sum_{j\neq i}r_{ij}^{-m}$, where $r_{ij}$ is the mean pairwise distance between the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ loci, and $m>0$ is an adjustable parameter. The centrality of a locus is higher when it is in proximity to other loci. Using $m=3$, we plotted $C_{i}$ against $M_{1}(t)$ at $t=10^{2}\mathrm{s}$ (Fig. 6(c)). The results show a negative correlation between the diffusivity of loci and the centrality measure. We recognize that the inverse of the pairwise distance $r_{ij}$ is correlated with the contact probability $p_{ij}$. We then inspect the correlation between diffusivity $M_{1}(t=10^{2}\mathrm{s})$ and $\sum_{j\neq i}p_{ij}$ and find that these two quantities are indeed anticorrelated (Fig. 6(d)).
@@
Figure 6: (a) Single-locus mean square displacement $M_{1}(t)$. Each curve corresponds to an individual locus. (b) Histogram of the diffusion exponent $\alpha$ and diffusion constant $D$. (c) Diffusivity, defined as $M_{1}(t)$ at $t=10^{2}\ \mathrm{s}$, as a function of closeness centrality $C_{i}$. (d) Diffusivity $M_{1}(t=10^{2}\mathrm{s})$ as a function of the sum of contact probabilities for each locus $i$.

### v3

@@ II Results
Next, we calculated the two-point mean square displacement $M_{2}(t)$ using Eq. 3 and the relation $M_{2}(t)=2\langle r_{ij}^{2}(t)\rangle-2G_{2}^{ij}(t)$. Use of Eq. 3 requires knowledge of the connectivity matrix $\bm{K}$, which is the byproduct of the determination of the 3D chromatin coordinates (Eq. 2). Fig. II(d) shows the time dependence of $M_{2}(t)$ predicted by our theory for the pairs of E-P distances. At long times, $M_{2}(t)$ saturates, approaching the different equilibrium values that depend on the given pair. The rate of approach depends on the specific enhancer and promoter pair.
@@
We then calculated the relaxation time $\tau$. Because the absolute value of the friction coefficient is unknown, we tuned it to achieve the best agreement between the theoretical prediction and experimental value for $\tau$. The fit parameter yields the unit length $l_{0}=147\mathrm{nm}$ and unit time is $\tau_{0}\approx 3.1\mathrm{s}$. Using $\tau_{0}=3.1$s and $l_{0}=140$nm, one finds $\xi=k_{\mathrm{B}}T\,\tau_{0}/l_{0}^{2}\approx 6.5\times 10^{-7}\,\mathrm{N\,s/m}$. Using Stokes’ law with $r=l_{0}/2$, we estimate the viscosity of the environment to be $\eta=\xi/(6\pi r)\approx 0.5\,$Pa·s. We used these values to calculate the theoretical predictions for $\tau$ versus genomic distance $s$ to compare with experiments. The theoretical prediction for the scaling exponent, $\gamma$ in $\tau\sim s^{\gamma}$, is $\approx 0.8$ and for experimental data is $\approx 0.7$ (Fig. II(e)). It is important to note that while $l_{0}$ and $\tau_{0}$ are adjustable parameters used to calculate the absolute values of $\tau$, they do not affect the scaling exponent $\gamma$. We consider the agreement for $\tau$, and especially $\gamma$, between experiment and theory to be striking because the only information that is used in the calculation is the Micro-C static contact map.
@@
Randomly Shuffled Sequence: We then wondered if the introduction of randomness in the Micro-C contact map would lead to a discrepancy between theory and experiment. To this end, we randomly shuffled the pairwise distances in the distance map but retained the first off-diagonal elements. In this way, the polymeric nature of the structure is preserved, but the specific WT (wild type) pattern in the Micro-C contact map is destroyed. We then applied HIPPS-DIMES on the shuffled distance map to obtain the connectivity matrix. Comparison of the contact maps between the WT and the randomly shuffled sequence shows (fig. S7(a)) that the pattern in the WT contact map is fully disrupted. fig. S7(b) shows the mean pairwise distances as a function of the genomic distance. For $s<10^{2}$ bps, the mean distance grows roughly as a power law with an exponent of 0.6, demonstrating the polymeric nature is preserved in the random shuffled system. At $s>10^{2}$ bps, the mean distances reach the plateau, which is a result of random shuffling. The relaxation time $\tau$ for the shuffled sequence is insensitive to the genomic distance (fig. S7(c)), which is consistent with the saturation of mean distances. The purpose of this calculation is to show that the scaling of $\tau$ as a function of $s$ cannot be captured in random heteropolymer. The chromatin sequence, reflecting the patterns of activity depicted in histone modifications, and the associated 3D structures should be accounted for precisely.
@@
Figure 2: Locus-specific relaxation times scale with spatial separation of chromatin loci in Drosophila embryo cells. (a) Heatmap comparison of the mean spatial distances (lower triangle) and relaxation time $\tau$ (upper triangle) across all locus pairs. (b) Log–log scatter plot of the relaxation time $\tau$ versus mean spatial distance $\langle r\rangle$. Black circles are experimental measurements; blue squares are HIPPS-DIMES predictions. The red dashed line is a power-law fit $\tau\sim\langle r\rangle^{2.7}$ highlighting the non-Rouse scaling.
@@
Loci-dependent relaxation times: Given the remarkable success of our theoretical approach in quantitatively explaining the experimental findings, we calculated all the pairwise relaxation time $\tau_{ij}$ where $i,j$ are the loci pair indices. The value of $\tau_{ij}$ depends on both $i$ and $j$, and not merely on the genomic distance $s=|i-j|$ as in the case of a homopolymer. On an average, the relaxation time correlates with both $s$ and the mean spatial distance $\langle r_{ij}\rangle$ in a non-trivial manner. Fig. 2(a) shows the mean spatial distance map and the relaxation time map, clearly establishing the correlation between the two quantities (see fig. S13 for zoom-in view of the one Mb scale comparison). Fig. 2(b) shows the scatter plot of $\langle r_{ij}\rangle$ versus $\tau_{ij}$. The results show that, on an average, they are related as $\tau_{ij}\sim\langle r_{ij}\rangle^{2.7}$, which differs substantially from the prediction for the Rouse and the FG models. In particular, the scaling exponent $\sim 2.7$ is substantially smaller than the Rouse model prediction ($=4$) and the value for FG ($=5$).
@@
We then wondered whether the observed scaling can be deduced by considering an effective homopolymer model, in which the mean distance matrix is calculated as the average of the wild-type (WT) distance map over fixed genomic distances. The calculation is intended to assess if a modified scaling relation could be used with the mean distance that is calculated from the contact maps. We computed $r(s)$ by averaging $\langle r_{ij}\rangle$ over $s=|j-i|$, and then applied HIPPS-DIMES to obtain the relaxation time. fig. S8 shows the calculated average distance map, demonstrating that $r(s)\sim s^{1/4}$ for $s>10^{2}\ \mathrm{kb}$. figs. S8(c) and 8(d) show that the relaxation time scales with genomic distance as $\tau\sim s^{1.1}$ and with mean pairwise distance as $\tau\sim r^{4}$. Both of these scaling relations are different from the results obtained by considering the full WT contact map. This further demonstrates that arguments in standard polymer physics for homopolymer do not hold for the chromatin, thus underscoring the importance of considering the complete information embedded in the WT contact map. Together, these results show that the relaxation process between a pair of chromatin loci is much faster than predicted by standard polymer models, which provides a structural basis for interpreting the key experimental finding [36]. Importantly, the rapid dynamics between distal loci can be explained by taking into account the actual 3D coordinates of the chromosomes.
@@
Finally, we examined whether enhancer-promoter (E-P) relaxation is faster than that of non-E-P pairs at the same genomic distances. For each of the six reporter separations (at 58 kb, 82 kb, 149 kb, etc.), we calculated the distribution of relaxation times for all loci pairs separated by the same genomic separation. The results, shown in fig. S9, reveal distinct trends depending on the on the genomic distance. For very short genomic distances (enhancer-promoter pairs 1 and 2, with separations of 58 kb and 82 kb), the mean relaxation times (133 s and 134 s, respectively) are substantially shorter than the mean relaxation times of all pairs at the same genomic distance (207 s and 240 s). In contrast, for larger genomic distances (all other enhancer-promoter pairs), the mean relaxation times are consistently longer than the corresponding mean values of all pairs with the same separation. Because these specific pairs are only available in the current experimental measurements, we cannot generalize further to other E–P combinations.
@@
Plausible mechanism for rapid relaxation dynamics in chromatin: To explore the underlying mechanism for the rapid relaxation times between chromatin loci, we calculated the spectrum of eigenvalues, $\lambda_{p}$, associated with, $\bm{K}$, the connectivity matrix. Interestingly, Fig. 3(a) shows that the scaling of $|\lambda_{p}|$ with $p$ has a complex structure. There are three distinct regimes in the variation of $\lambda_{p}$ with the normal mode $p$. For $p\leq 10$, we find that $|\lambda_{p}|\approx p^{1.2}$. In the second regime, $10\leq p\leq 50$ the eigenvalues increase as $|\lambda_{p}|\sim p^{3}$. Finally, $|\lambda_{p}|\sim p^{1.5}$ for $p\geq 50$. The complicated spectrum for chromatin should be contrasted with the Rouse model for which $|\lambda_{p}|\sim p^{2}$ where the inverse of $|\lambda_{p}|$ maybe interpreted as the relaxation time of $N/p$ segments of the chain. Notably, the smaller scaling exponents in the small $p$ regime (Fig. 3(a)) compared to the $p^{2}$ scaling of the Rouse model supports the finding that chromatin relaxes more rapidly, consistent with the results for relaxation time $\tau$. In general, for a homopolymer, we expect that ‘$|\lambda_{p}|\sim p^{x}$, which implies that the end-to-end relaxation times ($\tau_{ee}$) scale with respect to the chain length with the same exponent, $\tau_{ee}\sim N^{x}$. For instance, for the Rouse model, $|\lambda_{p}|\sim p^{2}$ and $\tau_{ee}\sim N^{2}$. If we assume that a similar power law relationship holds between $\tau_{ee}$ and $p$ in chromosomes, then expect that $\tau_{ee}\sim N^{1.2}$ if $s\gtrapprox$ 400 kb. To test this prediction, we calculated the end-to-end-distance relaxation times in chromosomes with different lengths. The HIPPS-DIMES-based calculation shows that the end-to-end relaxation time roughly scales linearly with $N$ (Fig. 3(b)) as $\tau_{ee}$ varies by over five orders of magnitude. We also computed the eigenvalues for the randomly shuffled system. fig. S10(a) shows that $|\lambda_{p}|$ becomes independent of $p$ for $p\leq 40$, which is is consistent with the findings that $\tau$ is insensitive to the genomic distance in the shuffled system. Together, these results suggest that the dynamics of chromosomes are dependent on the sequence and the length scale, which is reflected in the observation that the $|\lambda_{p}|$ exhibits three distinct scaling regimes at different $p$ (different length scales).
@@
Figure 3: Eigenvalue spectrum and chain-length dependence of the end-to-end relaxation times in Drosophila embryo cells. (a) Scaling of the eigvenvalues, $|\lambda_{p}|$, of the connectivity matrix as a function of mode index $p$, revealing three distinct scaling regimes. (b) End-to-end relaxation time $\tau_{ee}$ as a function of chain length $N$. The dashed line shows a power-law fit.
@@
First-passage time of contact between distal loci: A functionally relevant biophysical property related to the two-point relaxation time is the first-passage time of contact between a pair of chromatin loci. A simpler, well-studied, and instructive version of this problem in polymer physics is the cyclization process, which concerns the first-passage time for two ends of a polymer chain to meet [48, 49, 50, 46]. Let us denote the first-passage time of contact as $\tau_{c}$, which is determined by the search process by which two loci meet. It can be shown that $\tau_{c}$ is directly connected to the two-point relaxation and is governed by the relaxation dynamics between the loci [46, 50], assuming that the threshold for establishing contact is not small. Using a contact threshold of $r_{c}=l_{0}=147\mathrm{nm}$, we estimated the mean FPT, $\langle\tau_{c}\rangle$, for all pairs of the chromatin loci (see Supplementary Materials for method to estimate mean FPT). Fig. 4(a) shows that the domain along the diagonal in the $\tau_{c}$ map visually matches the contact domains in the contact map. We then calculated $\tau_{c}$ by averaging over the spatial distance $r$ with fixed genomic distance $s$. Fig. 4(b) shows that, on an average, the mean $\tau_{c}$ of contact between two chromatin loci scales as $\langle r\rangle^{3.4}$.
@@
We also compared our theoretical predictions with results computed using experimental loci trajectory data. The method for calculating the mean FPT from experimental trajectories is described in the Supplementary Materials. As shown in Fig. 4(b), the experimental results (represented by triangle symbols) align closely with our predictions, demonstrating excellent quantitative agreement between the experiment and the theory.
@@
Interestingly, the calculated scaling exponent of 3.4 is close to 3, which is the theoretical prediction by Szabo, Schulten, and Schulten (SSS) [49] who derived the first-passage time of contact under the assumption that two-point diffusion is governed by dynamics in a potential of mean force, which can be calculated analytically for the Rouse model. The potential of mean force is determined from the distribution of distances between two loci which further leads to [49, 46, 51],
@@
where $p(x)$ is the equilibrium distribution of distances $x$ between the two loci with the mean distance $\langle r\rangle$, $r_{c}$ is the threshold distance distance for contact, and $D$ is the effective diffusion constant. It can be shown using Eq. 4 that $\tau_{c,\mathrm{SSS}}\sim D^{-1}r_{c}^{-1}\langle r\rangle^{3}$. This result is consistent with the visual similarity between the contact map and the contact time map shown in Fig. 4(a), as the contact probability scales with the mean distance with an exponent of 3 (in the HIPPS-DIMES model). It is intriguing that although the SSS theory fails to predict the correct scaling in the Rouse model, it provides a better description of chromatin dynamics.
@@
The value of $\tau_{c}$ is important in describing the dynamics of enhancer-promoter (EP) communication and potentially in understanding EP-regulated gene expression. If we assume that gene expression is initiated by the formation of contact between promoter and enhancer, then the transcription rate can be expressed as $k=1/(\tau_{\mathrm{d}}+\tau_{c})$, where $\tau_{\mathrm{d}}$ is the time required for the downstream processes that ensue after the establishment of contact. By expressing $\tau_{c}$ as a function of the contact probability $p_{c}$, with $\tau_{c}\sim\tau_{0}p_{c}^{-\theta}$, we obtain $\hat{k}=k/k_{\mathrm{max}}=1/(1+(\tau_{0}/\tau_{d})p_{c}^{-\theta})$. This equation can be considered as the dynamic analog of the Hill equation with the cooperativity parameter $\theta$. Such an equation has been used to model the mean mRNA number as a function of the contact probability between promoter and enhancer [52].
@@
Figure 4: First-passage times for contact between chromatin loci in Drosophila embryo cells. (a) Comparison between heatmap for the mean first-passage time of contact between chromatin loci (upper triangle), $\langle\tau_{c}\rangle$, and the contact map (lower triangle). The contact is defined with a distance threshold of $r_{c}=147\mathrm{nm}$. (b) $\langle\tau_{c}\rangle$ scales as the mean distance $\langle r\rangle$ as a power-law. The dashed line is a guide for the eye. Triangle symbols are the data computed using experimental trajectory data [36] (See Supplementary Materials).
@@
Single-Locus Dynamics and Centrality Measure: Having demonstrated that our theory quantitatively reproduces the experimental data for the two-point dynamics in chromatin loci, we explore the predictions of single-locus dynamics. The single-locus mean square displacement, $M_{1}(t)$, is computed using Eq. 14 in the Supplementary Materials. The prediction for $M_{1}(t)$ is shown in Fig. 5(a), where each line represents a single chromatin locus. Between the time scale of $1\mathrm{\ s}<t<10^{5}\mathrm{\ s}$, $M_{1}(t)$ scales approximately with an exponent of $\sim 0.5$, which is close to the prediction of the Rouse model. Next, we calculated the diffusion exponent $\alpha$ and diffusion constant $D$ for each locus by fitting $M_{1}(t)$ in the time range $1\mathrm{s}<t<10^{5}\mathrm{s}$ using $M_{1}(t)=Dt^{\alpha}$. The histogram of $\alpha$ and $D$ reported in Fig. 5(b), yields the loci averages of $\langle\alpha\rangle=0.52$ and $\langle D\rangle=0.024\ \mu\mathrm{m^{2}s^{-1/2}}$. Fig. 5(b) shows broad distributions of both the exponent $\alpha$ and the effective diffusion coefficient, indicating that single-locus diffusion is heterogeneous.
@@
We then investigated the factors that determine the variance in single-locus diffusion. We hypothesize that chromatin loci should generally diffuse more slowly if they have higher local density. Inspired by concepts in graph theory [53], we defined the closeness centrality measure based on the mean pairwise distances. Let us define the centrality of a single locus as $C_{i}$, where $i$ is the locus index, as, $C_{i}=\sum_{j\neq i}\langle r_{ij}\rangle^{-m}$, where $\langle r_{ij}\rangle$ is the mean pairwise distance between the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ loci, and $m>0$ is an adjustable parameter. The centrality of a locus is higher when it is in proximity to other loci. Using $m=3$, we plotted $C_{i}$ against $M_{1}(t)$ at $t=10^{2}\mathrm{s}$ (Fig. 5(c)). The results show a negative correlation between the diffusivity of loci and the centrality measure. We recognize that the inverse of the pairwise distance $r_{ij}$ is correlated with the contact probability $p_{ij}$. We then inspect the correlation between diffusivity $M_{1}(t=10^{2}\ \mathrm{s})$ and $\sum_{j\neq i}p_{ij}$ and find that these two quantities are indeed anticorrelated (Fig. 5(d)).
@@
Figure 5: Single-locus dynamics and their dependence on chromatin network centrality in Drosophila embryo cells. (a) Single-locus mean square displacement $M_{1}(t)$. Each curve corresponds to an individual locus. (b) Histogram of the fitted diffusion exponent $\alpha$ and diffusion coefficients $D$. (c) Locus-specific diffusivity, defined as $M_{1}(t=10^{2}\ \mathrm{s})$, versus closeness centrality $C_{i}=\sum_{j\neq i}\langle r_{ij}\rangle^{-m}$ with $m=3$. (d) Scatter plot of locus-specific diffusivity, defined as $M_{1}(t)$ at $t=10^{2}$ s, versus total contact connectivity $\sum_{j\neq i}p_{ij}$ as a function of the sum of contact probabilities for each locus $i$.